∫[cosx]dx limit is from 0 to pi/2. [] is integral part of x.

∫[cosx]dx
limit is from 0 to pi/2. [] is integral part of x.

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int_{0}^{\pi /2}[cosx]dx
0 < x < \frac{\pi }{2}
[cosx] = 0
I = \int_{0}^{\pi /2}(0)dx
I = 0